30. The skier with mass M = 40 kg moves with speed V = 2m / s and throws a ball with mass m = 3 kg in the direction of movement a) Find how much should be the speed of the ball so that after throwing the skier stops b ) With the help of calculations check whether or not the skier can realize this situation, when it is known that he is able to communicate with the ball at a speed not exceeding 40 km / h.


1
Expert's answer
2021-02-05T11:48:07-0500

a) To find how much the speed of the ball should be so that after throwing the skier stops, we need to apply law of momentum conservation: the final momentum of the skier must be 0:


"(M+m)\cdot v=M\cdot0+mu,\\\space\\
u=v\frac{M+m}{m}=28.67\text{ m/s}."

Convert this to km/h:


"u_\text{km/h}=u\cdot3600/1000=103.2\text{ km/h}."

Thus, he can't stop himself by throwing this snowball.